Suppose there is a 60 percent chance it will rain today, a 50 percent chance it ll rain tomorrow, and a 30 percent chance it will not rain on either day. What is the probability that it will rain both today and tomorrow. Choose the closest answer. a. 0.0 b. 0.1 С. 0.2 d. 0.3 e. 0.4 f, 0.5 g, 0.6 h, 0.7
Answer #1

P(it will rain on either day) = 1 - P(it will not rain on either day) = 1 - 0.3 = 0.7

P(it will rain both today and tomorrow) = 0.6 + 0.5 - 0.7 = 0.4

Option - e is correct.